The heating coil rated 1kw is used to evaporated 700g of water at 100 degree.if specific latent heat of vaporization of water is 2.e×10^6, the time taken to evaporate all the water is?
1
Expert's answer
2019-02-04T10:01:48-0500

Let's write the formula for the amount of heat required to change the state of water at

"100 \u2103"

to a gas:

"Q = mL,"

here,

"Q"

is the amount of heat required to change the state of water at

"100 \u2103"

to a gas,

"m = 0.7 kg"

is the mass of the water and

"L = 2 \cdot 10^6 \frac{J}{kg}"

is the specific latent heat of vaporization of water.

From the other hand, we can write the formula for the amount of heat that a coil provides at the rate of

"1 kW"

during time

"t"

:

"Q = Pt,"

here,

"t"

is the time taken to evaporate all the water.

Then, we can equate both expressions and find the time taken to evaporate all the water:

"mL = Pt,""t = \frac{mL}{P} = \frac{0.7 kg \cdot 2 \cdot 10^6 \frac{J}{kg}}{1000 W} = 1400 s."

Answer:

"t = 1400 s"

.